OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (3,-4)? Find the equation, in standard form, of the line passing through the points (3,-4) and (5,1). A. 5x+2y=23 B. y= 5/2x -23/2 C. 5x-2y=23 D. 2x-3y=9 is it B?

jimthompson5910 (jim_thompson5910):

No, choice B is NOT in standard form (it's in slope-intercept form).

OpenStudy (anonymous):

C. 5x - 2y = 23

OpenStudy (campbell_st):

find the gradient of the line rise/run m = ( 1 - -4)/(5 - 3) m = 5/2 the line is in the form y = 5/2 x + b substitute either point, say x = 3 and y - 4 to find the value of b, the y-intercept.

OpenStudy (anonymous):

Thanks for the medal! :)